The price of Swearengen, Inc., stock will be either $75 or $97 at the end of the year. Call options are available with one year to expiration. T-bills currently yield 7 percent.

  

a.

Suppose the current price of the company's stock is $86. What is the value of the call option if the exercise price is $71 per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Call value $   

  

b.

Suppose the current price of the company's stock is $86. What is the value of the call option if the exercise price is $81 per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Call value $   
Answer #1

a

Upmove (U)= High price/current price=97/86=1.1279
Down move (D)= Low price/current price=75/86=0.8721
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.07*1)-0.8721)/(1.1279-0.8721)=0.78344
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(97-71,0)
=Max(26,0)
=26
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(75-71,0)
=Max(4,0)
=4
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.07*1)*(0.783441*26+(1-0.783441)*4)
=19.8

b

Upmove (U)= High price/current price=97/86=1.1279
Down move (D)= Low price/current price=75/86=0.8721
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.07*1)-0.8721)/(1.1279-0.8721)=0.78344
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(97-81,0)
=Max(16,0)
=16
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(75-81,0)
=Max(-6,0)
=0
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.07*1)*(0.783441*16+(1-0.783441)*0)
=11.69